# Scalable and Ethical Insider Threat Detection through Data Synthesis and Analysis by LLMs

### v1

@@ II-A LLM Selection
Research was performed on LLMs that met specific use criteria that would be suitable for review generation and sentiment analysis. The most important LLM criteria were natural language processing with native sentiment analysis capabilities, an API for interface with a python script to pass data from a comma separated value (CSV) file to the LLM, and the ability write results back to a CSV file. For sentiment-related tasks, Sonnet 3.5 and GPT-4o are outstanding [weitl-harms_using_2024], while demonstrating that they can capably produce synthetic sentiment data [hastings_utilizing_2024]. For the tailored prompts described in this section, prompts from [hastings_utilizing_2024] were used as a starting point.

### v2

@@ II-A LLM Selection
Research was performed on LLMs that met specific use criteria for review generation and sentiment analysis. Critical LLM criteria were natural language processing with native sentiment analysis capabilities, an API that could interface with Python to pass data from a comma separated value (CSV) file to the LLM, and the ability to write results back to a CSV file. For sentiment-related tasks, Sonnet 3.5 and GPT-4o are outstanding [weitl-harms_using_2024], while demonstrating that they can capably produce synthetic sentiment data [hastings_utilizing_2024]. Both the tailored prompts described in this section, and the Python code for accessing the API, were adapted from [hastings_utilizing_2024]. The code was adjusted for the current research by Claude and manually verified, while the prompts were adapted manually.